OpenStudy (anonymous):

find the domain of the function: f(x)= 2x

OpenStudy (anonymous):

domain is all real numbers because you can multiply any number by 2

OpenStudy (anonymous):

how would i write that out? like if i was working it out?

OpenStudy (amistre64):

in general, the domain is the set of numbers that does NOT make the function go bad

OpenStudy (anonymous):

first of all the domain is supposed to be part of the definition of the function, but in this case the question really is "what is the largest set of numbers for which you can evaluate the function". in general you have to worry about two things: you cannot divide by 0 and you cannot take the square root of a negative number

OpenStudy (anonymous):

so if you see a polynomial like \[f(x)=2x\] or \[g(x)=x^2-x\] just say "all real numbers'

OpenStudy (anonymous):

if you see \[f(x)=\frac{1}{x-1}\] you would say "all real numbers except 1"

OpenStudy (anonymous):

and if you see \[f(x)=\sqrt{x+2}\] you would say \[x\geq -2\]

OpenStudy (anonymous):

I just need to know how to work out f(x)=2x because i dont't understand it